# 2d Conformal Field Theories on Magic Triangle

### v1

@@ 2.2 Modular linear differential equation
In this section, we briefly review modular linear differential equations (MLDEs) and their fermionic version, namely MLDEs associated with congruence subgroups of level two. There exist several equivalent formulations of MLDEs, including those based on ordinary derivatives, Serre derivatives, and Rankin–Cohen brackets; see, for example, Nagatomo et al. (2024). For our purposes, it is most convenient to work with the Serre derivative
@@
where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for $SL(2,\mathbb{Z})$, generated by the Eisenstein series $E_{4}$ and $E_{6}$. The non-negative integer $l$ is referred to as the index of the MLDE. When $l=0$, the equation is called holomorphic or monic; otherwise it is meromorphic.
@@
When these solutions are identified with the characters of a 2d RCFT, the exponents are related to the central charge $c$ and conformal weights $h_{i}$ by
@@
One of the important applications of MLDE is to classify 2d RCFTs with given number of characters, or bootstrap new 2d RCFTs. This program requires a classification of CFT-type solutions for given pairs $(d,l)$. In the next subsection, we review the classical work of Mathur, Mukhi, and Sen on MLDEs with $(d,l)=(2,0)$ Mathur et al. (1988). More recent progress on the classification of CFT-type MLDE solutions can be found, for example, in Kaidi et al. (2021).
It is natural to generalize the properties of the characters of 2d RCFTs to the case of 2d fermionic RCFTs. In the presence of fermions, one must distinguish between periodic (Ramond) and anti-periodic (Neveu–Schwarz) boundary conditions along a circle, leading to four sectors depending on the boundary conditions along the two cycles of the torus. Accordingly, fermionic RCFTs admit four types of characters: NS, $\widetilde{\rm NS}$, R, and $\widetilde{\rm R}$. These transform as vector-valued modular forms for the congruence subgroups $\Gamma_{\theta}(2)$, $\Gamma^{0}(2)$, $\Gamma_{0}(2)$, and $SL(2,\mathbb{Z})$, respectively; see, for instance, Bae et al. (2021a); Duan et al. (2023). The fermionic characters of a fermionic RCFT satisfy corresponding fermionic MLDEs. For example, the Neveu–Schwarz characters obey an MLDE of the form (8), where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for the congruence subgroup $\Gamma_{\theta}(2)$.
@@ 2.3 Mathur–Mukhi–Sen’s discovery
This equation is also known in the literature as the Kaneko–Zagier equation Kaneko and Zagier (1998). Mathur, Mukhi, and Sen showed that only ten discrete values of the parameter $\lambda$ admit CFT-type solutions. Moreover, for each such value of $\lambda$, the corresponding solutions coincide precisely with the characters of level-one Wess–Zumino–Witten models associated with the extended Cvitanović–Deligne exceptional series:
@@
We refer to the sequence (11) as the Mathur–Mukhi–Sen series, and denote them as $\mathcal{T}^{(\nu,5)}$ or $\mathcal{T}^{(5,\nu)}$ theories with $\nu$ defined earlier associated to the CD exceptional series. This notation will be consistent with magic triangle at level one in Table 2. The associated CFT data are summarized in Table 3. Here $\mathrm{Vir}_{5,2}^{\rm eff}$ denotes the effective Lee–Yang model with effective central charge $2/5$. The entry $(E_{7+1/2})_{1}$ corresponds to the solution of the MLDE with central charge $c=38/5$. Mathur, Mukhi, and Sen observed that, if interpreted as a level-one WZW model, the corresponding symmetry algebra would have dimension $190$ and dual Coxeter number $24$, and hence could not arise from a semisimple Lie algebra. It was subsequently shown that these solutions can be realized as the characters of an intermediate vertex subalgebra of $(E_{8})_{1}$ Kawasetsu (2014). An alternative interpretation identifies them as the characters of a minimal $\mathcal{W}$-algebra with central charge $32/5$ Kawasetsu (2018). The $(E_{8})_{1}$ theory has the single vacuum character $j(\tau)^{1/3}$ Mathur et al. (1988). In this case, the second order MLDE has a spurious solution, see the Appendix C.
@@
Table 3: 2d CFTs associated to the extended CD exceptional series at level 11.
Recall that the Wess–Zumino–Witten model $(\mathfrak{g})k$ has central charge
$c_{k}=\frac{kd_{\mathfrak{g}}}{k+h^{\vee}}.$ (12)
For the extended Cvitanovi’c–Deligne exceptional series, both the dimension $d_{\mathfrak{g}}$ and the dual Coxeter number $h^{\vee}$ can be expressed as functions of the parameter $\nu$. At level one, the corresponding central charge and the conformal weight of the non-vacuum primary field are given by
@@
In terms of the parameter $\nu$, the associated second-order holomorphic MLDE can be written compactly as $\Delta_{\nu}\chi=0$, where $\Delta_{\nu}\chi=0$, where
@@
where $\mathcal{N}$ is natural integer, capturing the degeneracy of the non-vaccum primary operator. The characters of the Mathur–Mukhi–Sen series admit the expansions
$\chi_{0}$ $=q^{-c/24}(1+d_{\mathfrak{g}}q+\cdots),$
$\chi_{1}$ $=q^{-c/24+h}(R+\cdots),$ (16)
where $R$ denotes the dimension of the fundamental representation of $\mathfrak{g}$. The $S$-matrix arises from the modular transformation of the characters,
@@
As a $2\times 2$ matrix, the $S$-matrix satisfies $\det S_{(\nu)}=-1$ and
@@
It is well known that for the CD exceptional series at level $1$, the RCFTs pair together as a coset with respect to $\mathcal{T}^{(5,5)}=(E_{8})_{1}$ Mathur et al. (1988). These cosets are
@@
$\mathcal{T}^{(5,\nu)}=\frac{(E_{8})_{1}}{\mathcal{T}^{(1/\nu,5)}}.$ (28)
The non-vacuum conformal weight $h_{\nu}$ of $\mathcal{T}_{5,\nu}$ satisfies the following relation $h_{\nu}+h_{1/\nu}=1.$ The degeneracy $\mathcal{N}_{(\nu)}$ and $S$-matrix $S_{(\nu)}$ satisfy
$\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)},\ \ S_{(\nu)}=S_{(1/\nu)}.$ (29)
@@
$\chi_{0}^{(\nu)}\chi_{0}^{(1/\nu)}+\mathcal{N}_{(\nu)}\chi_{1}^{(\nu)}\chi_{1}^{(1/\nu)}=j^{1/3},$ (30)
which is $S$-invariant. The above bilinear relation in the subleading $q$-order implies
@@
Multiplying $\mathcal{N}_{(\nu)}R_{(1/\nu)}^{2}$ one can also obtain the relation (31).
The characters of WZW models for CD exceptional series at level one have many other interesting properties such as Hecke relations Harvey and Wu (2018). Besides, it is recently shown that the vacuum character of WZW models for CD exceptional series at level 1 can be constructed from the monodromy trace of some Argyres–Douglas theories Kim and Song (2024).
@@ 3 Magic triangle at level one
A principal motivation of this work is to extend the discovery of Mathur, Mukhi, and Sen concerning second-order holomorphic MLDEs and the Cvitanovi’c–Deligne exceptional series to the full magic triangle. To this end, we begin by analyzing the level-one Wess–Zumino–Witten models associated with the ordinary Lie groups appearing in the original magic triangle of Table 1. For each resulting rational conformal field theory, we compute the characters and show that they satisfy a uniform two-parameter family of fourth-order holomorphic MLDEs.
Assuming that this MLDE framework extends to the enlarged magic triangle, we solve for the characters corresponding to the previously unknown entries and identify the associated 2d RCFTs. The complete set of resulting theories is summarized in Table 2, which we interpret as the extended magic triangle at level one. We further uncover several structural properties of this level-one triangle, including the magic coset, magic bilinear relation and matrix-valued modularity of the characters.
@@ 3.1 The fourth order MLDE
Let us first consider the level-one theories for the original magic triangle in Table 1. Substituting the dimension and dual Coxeter number formulas (5), (6) into the central charge formula of WZW model (12), we obtain the central charge for the level-one theory $\mathcal{T}^{(\mu,\nu)}$ as
@@
The exponents $\alpha_{i}=-c/24+h_{i}$ in the corresponding characters $\chi_{i}$ take values,
@@
Denote the characters of the theory $\mathcal{T}^{(\mu,\nu)}$ as
@@
where $R_{i}$ are the dimensions of the irreducible representation of $\mathfrak{g}^{(\mu,\nu)}$. The coefficients $a_{n},b_{n},c_{n},d_{n}$ are the sum of the dimensions of the irreducible representations for the Lie algebra. The modular invariant partition function of $\mathcal{T}^{(\mu,\nu)}$ on torus is
@@
where $\mathcal{M}_{i}$ are natural numbers that capture the degeneracy of primaries fields.
The first main result of this paper is as follows. We find that the characters $\chi_{i}$ of the theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the following uniform fourth-order modular linear differential equation
@@
where the three coefficients $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ are characterized by $(\mu,\nu)$ as follows:
@@
Taking $\nu=5$, the 4th order MLDE factorizes to
@@
The differential operator can be written as $\Delta_{1/\mu}\circ\Delta_{\mu}$ where $\Delta_{\mu}$ is the second order operator given in (14). This factorization demonstrates that the fourth-order MLDE (41) provides a natural generalization of the second-order Mathur–Mukhi–Sen equation.
@@
It is natural to postulate that the above expressions for the central charge, conformal weights, and the modular linear differential equation (41) continue to hold for the extended magic triangle. This assumption allows us to identify two-dimensional rational conformal field theories beyond the original magic triangle of ordinary Lie groups. Indeed, upon specializing to $\nu=4$, the equation (41) reduces to a one-parameter family of fourth-order MLDEs governing minimal $\mathcal{W}$-algebras Kawasetsu and Sakai (2018). In this case, the characters of the theory $\mathcal{T}^{(\mu,4)}$ coincide with the characters of Ramond-twisted irreducible modules of the minimal $\mathcal{W}$-algebra $\mathcal{W}{-h^{\vee}/6}(\mathfrak{g},f\theta)$, where $\mathfrak{g}$ belongs to the Cvitanovi’c–Deligne exceptional series. More generally, by explicitly solving the MLDE (41) and interpreting its solutions as RCFT characters, we are led to identify a class of theories $\mathcal{T}^{(\mu,\nu)}$ extending those associated with the original magic triangle. [^1] Note that two theories ${\mathcal{T}}^{(\frac{2}{3},2)}$ and ${\mathcal{T}}^{(1,\frac{3}{2})}$ have different central charges $4/5$ and $6/5$ respectively, even though they have the same dimensions 0 and dual Coxeter numbers 1-1.
@@
  • $\mathcal{T}^{(2,1)}$ is a unitary RCFT with $c=2$ and conformal weights $h_{i}=0,(\frac{1}{6})_{6},(\frac{1}{2})_{3},(\frac{2}{3})_{2}$. The VOA was constructed in Kitazume et al. (2002), see Remark 5.8 there for all its irreducible modules. The RCFT is denoted as $M_{2}$ in (Duan et al., 2022, Section 5.6). The four characters have appeared as CFT-type solutions of fourth order holomorphic MLDE in (Kaidi et al., 2021, Table 6). The $q$-series of the characters can be found in (Duan et al., 2022, Equation (5.34)). The $M_{2}$ can be realized by the coset construction $(D_{4})_{1}/(A_{2})_{1}$ (Duan et al., 2022, Section 5.6). We notice that it can also be realized as a non-diagonal modular invariant of $(U_{1})_{2}\times(U_{1})_{6}$, which is consistent with magic triangle element $U_{1}^{2}$.
@@
  • $\mathcal{T}^{(3/2,1)}=D_{2\rm A}$ is a unitary RCFT associated to the 2A conjugacy class of the Monster group. It has $c=6/5$ and conformal weights $h_{i}=0,(\frac{1}{10})_{3},(\frac{1}{2})_{3},\frac{3}{5}$. The VOA can be constructed using a code in $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and as an extension of the product of Ising model $\mathrm{Vir}_{4,3}$ and tri-critical Ising model $\mathrm{Vir}_{5,4}$ Lam and Yamada (2000). For 2d RCFTs associated to various conjugacy classes and sporadic groups, see a good review in Bae et al. (2021b).
@@
  • $\mathcal{T}^{(3,2/3)}=\mathrm{Vir}_{6,5}^{\rm e}$ is a non-diagonal modular invariant of minimal model $\mathrm{Vir}_{6,5}$ which describes the 3-state Potts model. The VOA can be constructed using a code in $\mathbb{Z}_{3}$ Miyamoto (2001).
@@
  • $\mathcal{T}^{(3/2,3/2)}=(A_{1})_{8}^{\rm e}$ is the D-type modular invariant of $(A_{1})_{8}$. The VOA can be constructed by simple current extension.
@@
All RCFTs associated to the extended magic triangle are presented in Table 2. The data of the extended magic triangle and the RCFTs are collected in the Tables 5, 6, 7 in the Appendix A.
@@
Now we discuss the symmetry of the magic triangle at level one. It is easy to verify that under the exchange of $\mu$ and $\nu$, the central charge $c$, the conformal weight $h_{1}$ and the fourth order MLDE are invariant, while $h_{2},h_{3}$ are exchanged. This suggests that the $\mathcal{T}^{(\mu,\nu)}$ and $\mathcal{T}^{(\nu,\mu)}$ theories are isomorphic to each other, with only the labeling of $\chi_{2}$ and $\chi_{3}$ is exchanged. Therefore we only need to present half of the triangle in the Tables 5, 6, 7. In total, there are 30 non-equivalent RCFTs in the extended magic triangle at level one. The 16 of them come from ordinary WZW model of semisimple Lie groups, 8 of them come from intermediate CFTs, three of them from Virasoro minimal models, two of them from $U(1)$ theory, i.e. compact bosons on a circle, and the last one $D_{2\rm A}$ associated to Monster group. For all 30 theories on the magic triangle, the Fourier coefficients of the characters remain non-negative integers.
A few remarks are in order:

### v2

@@ 2.2 Modular linear differential equation
In this section, we briefly review modular linear differential equations (MLDEs) and their fermionic counter parts, namely MLDEs associated with congruence subgroups of level two. There exist several equivalent formulations of MLDEs, including those based on ordinary derivatives, Serre derivatives, and Rankin–Cohen brackets; see, for example, Nagatomo et al. (2024). For our purposes, it is most convenient to work with the Serre derivative
@@
where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for $SL(2,\mathbb{Z})$, generated by the Eisenstein series $E_{4}$ and $E_{6}$. The non-negative integer $l$ is referred to as the index of the MLDE. When $l=0$, the equation is called holomorphic or monic; otherwise meromorphic.
@@
When these solutions are identified with the characters of a 2d RCFT, the exponents are related to the central charge $c$ and conformal weights $h_{i}$ of primary fields by
@@
One of the important applications of MLDEs is the classification of 2d RCFTs with a given number of characters, as well as the bootstrapping of new 2d RCFTs. This program requires a classification of CFT-type solutions of MLDEs for given pairs $(d,l)$. In the next subsection, we review the classical work of Mathur, Mukhi, and Sen on MLDEs with $(d,l)=(2,0)$ Mathur et al. (1988). More recent progress on the classification of CFT-type MLDE solutions can be found, for example, in Kaidi et al. (2021).
It is natural to generalize the properties of the characters of 2d RCFTs to the case of 2d fermionic RCFTs. In the presence of fermions, one must distinguish between periodic (Ramond) and antiperiodic (Neveu–Schwarz) boundary conditions along a circle, leading to four sectors depending on the boundary conditions along the two cycles of the torus. In summary, fermionic RCFTs admit four types of characters: NS, $\widetilde{\rm NS}$, R, and $\widetilde{\rm R}$. They transform as vector-valued modular forms for the congruence subgroups $\Gamma_{\theta}(2)$, $\Gamma^{0}(2)$, $\Gamma_{0}(2)$, and $SL(2,\mathbb{Z})$, respectively; see, for instance, Bae et al. (2021a); Duan et al. (2023) for details. The fermionic characters of a fermionic RCFT satisfy corresponding fermionic MLDEs. For example, the NS characters obey an MLDE of the form (8), where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for the congruence subgroup $\Gamma_{\theta}(2)$. Then the fermionic MLDEs for $\widetilde{\rm NS}$ characters and the one for R characters can be obtained by modular transformation from the one for NS characters.
@@ 2.3 Mathur–Mukhi–Sen’s discovery
This equation is also known in the literature as the Kaneko–Zagier equation Kaneko and Zagier (1998). Mathur, Mukhi, and Sen showed that only for ten discrete values of the parameter $\lambda$ does the MLDE (10) admit CFT-type solutions satisfying with the unique vacuum condition. Moreover, for each such value of $\lambda$, the corresponding solutions coincide precisely with the characters of level-one Wess–Zumino–Witten models associated with the extended Cvitanović–Deligne exceptional series:
@@
We refer to the sequence (11) as the Mathur–Mukhi–Sen series, and denote these theories as $\mathcal{T}^{(\nu,5)}$ or $\mathcal{T}^{(5,\nu)}$, where $\nu$ is defined earlier and is associated with the CD exceptional series. This notation will be consistent with magic triangle at level one in Table 2. The associated CFT data are summarized in Table 3. Here $\mathrm{Vir}_{5,2}^{\rm eff}$ denotes the effective Lee–Yang model with effective central charge 2/52/5.333This is to exchange the labeling of the vacuum and non-vacuum operator of the original Lee–Yang model with c=22/5.c=-{22}/{5}. The entry $(E_{7+1/2})_{1}$ corresponds to the solution of the MLDE with central charge $c=38/5$. Mathur, Mukhi, and Sen observed that, if this solution were interpreted as a level-one WZW model, the corresponding symmetry algebra would have dimension $190$ and dual Coxeter number $24$, and hence could not arise from a semisimple Lie algebra. It was subsequently shown that these solutions can be realized as the characters of an intermediate vertex subalgebra of $(E_{8})_{1}$ Kawasetsu (2014). Accordingly, $(E_{7+1/2})_{1}$ is often reffered to as an intermediate VOA in the physics literature. An alternative interpretation identifies the solutions as the characters of Ramond-twisted modules of a minimal $\mathcal{W}$-algebra of type $E_{8}$ with central charge $32/5$ Kawasetsu (2018). Finally, the $(E_{8})_{1}$ theory has the single character, giveb by $j(\tau)^{1/3}$ Mathur et al. (1988). In this case, the second order MLDE admits a spurious solution, see Appendix C.
@@
Table 3: 2d CFTs associated to the extended Cvitanović–Deligne exceptional series at level 11.
Recall that the Wess–Zumino–Witten model $(\mathfrak{g})_{k}$ has central charge and conformal weights given by
$c_{k}=\frac{kd_{\mathfrak{g}}}{k+h^{\vee}},\qquad h_{\lambda}=\frac{C_{2}(R_{\lambda})}{2(k+h^{\vee})},$ (12)
where $\lambda$ is the weight and $C_{2}(R_{\lambda})$ is the quadratic Casimir invariant of representation $R_{\lambda}$. For the extended Cvitanović–Deligne exceptional series, both the dimension $d_{\mathfrak{g}}$ and the dual Coxeter number $h^{\vee}$ can be expressed as functions of the parameter $\nu$; the same is true for the quadratic Casimir invariant of the fundamental representation. At level one, the corresponding central charge and the conformal weight of the non-vacuum primary field are given by
@@
In terms of the parameter $\nu$, the associated second-order holomorphic MLDE can be written compactly as $\Delta_{\nu}\chi=0$, where
@@
where $\mathcal{N}$ is a natural number, capturing the degeneracy of the non-vacuum primary operator. The characters of the Mathur–Mukhi–Sen series admit the Fourier expansions
$\chi_{0}$ $=q^{-\frac{c}{24}}(1+d_{\mathfrak{g}}q+\cdots),$
$\chi_{1}$ $=q^{-\frac{c}{24}+h}(R+\cdots),$ (16)
where $R$ denotes the dimension of the fundamental representation of $\mathfrak{g}$. The degeneracy $\mathcal{N}$ and dimension $R$ for various $\nu$ are collected in Table 3. The $S$-matrix arises from the modular transformation of the characters,
@@
As a $2\times 2$ matrix not necessarily symmetric, the $S$-matrix satisfies $\det S_{(\nu)}=-1$ and
@@
It is well known that for the CD exceptional series at level $1$, the RCFTs pair together as a coset with respect to $\mathcal{T}^{(5,5)}=(E_{8})_{1}$ Mathur et al. (1988). These coset relations are
@@
$\mathcal{T}^{(5,\nu)}=\frac{\mathcal{T}^{(5,5)}}{\mathcal{T}^{(1/\nu,5)}}.$ (28)
The non-vacuum conformal weight $h_{\nu}$ of $\mathcal{T}_{5,\nu}$ theory satisfies the relation $h_{\nu}+h_{1/\nu}=1.$ The degeneracy $\mathcal{N}_{(\nu)}$ and $S$-matrix $S_{(\nu)}$ satisfy
$\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)},\qquad S_{(\nu)}=S_{(1/\nu)}.$ (29)
@@
$\chi_{0}^{(\nu)}\chi_{0}^{(1/\nu)}+\mathcal{N}_{(\nu)}\chi_{1}^{(\nu)}\chi_{1}^{(1/\nu)}=j^{1/3}.$ (30)
The above bilinear relation of characters in the subleading $q$-order implies
@@
By multiplying $\mathcal{N}_{(\nu)}R_{(1/\nu)}^{2}$ one can also obtain the relation (31).
The characters of WZW models for the CD exceptional series at level one possess many other interesting properties such as Hecke relations Harvey and Wu (2018). Besides, it is recently shown that the vacuum character of WZW models for CD exceptional series at level one can be constructed from the monodromy trace of some Argyres–Douglas theories Kim and Song (2024).
@@ 3 Magic triangle at level one
A principal motivation of this work is to extend the discovery of Mathur, Mukhi, and Sen concerning second-order holomorphic MLDEs and the Cvitanović–Deligne exceptional series to the full magic triangle. To this end, we begin by analyzing the level-one Wess–Zumino–Witten models associated with the ordinary Lie groups appearing in the original magic triangle of Table 1. For each resulting rational conformal field theory, we compute the characters and find that they satisfy a uniform two-parameter family of fourth-order holomorphic MLDE (41). Assuming that this MLDE framework extends to the enlarged magic triangle, we determine the characters corresponding to the previously unknown entries and identify the associated 2d RCFTs. The complete set of resulting theories is summarized in Table 2, which we interpret as the extended magic triangle at level one. We further uncover several elegant properties of this level-one triangle, including the magic coset structure, matrix-valued modularity of the characters with explicit S-matrices, the dimensions and degeneracies of primary operators, and the existence of five atomic models that together constitute all theories in the triangle.
@@ 3.1 The fourth order MLDE
Let us first consider the level-one theories for the original magic triangle in Table 1. Substituting the dimension and dual Coxeter number formulas (5), (6) into the central charge formula of WZW model (12), we obtain the central charge of the level-one theory $\mathcal{T}^{(\mu,\nu)}$ as
@@
The exponents $\alpha_{i}=-c/24+h_{i}$, appearing in the corresponding characters $\chi_{i}$, take values
@@
We denote the four characters of the theory $\mathcal{T}^{(\mu,\nu)}$ by
@@
where $R_{i}$ denote the dimensions of irreducible representations of $\mathfrak{g}^{(\mu,\nu)}$. The coefficients $a_{n},b_{n},c_{n}$, and $d_{n}$ are the sum of the dimensions of the irreducible representations for the Lie algebra. The modular invariant partition function of $\mathcal{T}^{(\mu,\nu)}$ on torus is
@@
where $\mathcal{M}_{i}$ are positive integers that capture the degeneracies of primary fields.
The first main result of this paper is the following. We find that the characters $\chi_{i}$ of the theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the uniform fourth-order modular linear differential equation
@@
where the three coefficients $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ are determined by $(\mu,\nu)$ as follows:
@@
Specializing to $\nu=5$, the 4th order MLDE factorizes as
@@
This differential operator can thus be written as $\Delta_{1/\mu}\circ\Delta_{\mu}$ where $\Delta_{\mu}$ is the second order operator given in (14). This factorization demonstrates that the fourth-order MLDE (41) provides a natural generalization of the second-order Mathur–Mukhi–Sen equation.
@@
It is natural to postulate that the above expressions for the central charge, conformal weights, and the modular linear differential equation (41) continue to hold for the extended magic triangle. This assumption allows us to identify two-dimensional rational conformal field theories beyond the original magic triangle of ordinary Lie groups. Indeed, upon specializing to $\nu=4$, we find the equation (41) reduces to a one-parameter family of fourth-order MLDEs governing some minimal $\mathcal{W}$-algebras Kawasetsu and Sakai (2018). In this case, the characters of the theory $\mathcal{T}^{(\mu,4)}$ coincide with the characters of Ramond-twisted irreducible modules of the minimal $\mathcal{W}$-algebra $\mathcal{W}_{-h^{\vee}/6}(\mathfrak{g},f_{\theta})$, where $\mathfrak{g}$ belongs to the Cvitanović–Deligne exceptional series. More generally, by explicitly solving the MLDE (41) and interpreting its solutions as RCFT characters, we are led to identify a class of theories $\mathcal{T}^{(\mu,\nu)}$ extending those associated with the original magic triangle. [^4] Note that two theories ${\mathcal{T}}^{(\frac{2}{3},2)}$ and ${\mathcal{T}}^{(1,\frac{3}{2})}$ have different central charges $4/5$ and $6/5$ respectively, even though they have the same dimensions 0 and dual Coxeter numbers 1-1.
@@
  • $\mathcal{T}^{(2,1)}$ is a unitary RCFT with $c=2$ and conformal weights $h_{i}=0,(\frac{1}{6})_{6},(\frac{1}{2})_{3},(\frac{2}{3})_{2}$. The VOA was constructed in Kitazume et al. (2002), see Remark 5.8 there for all its irreducible modules. This RCFT is denoted as $M_{2}$ in (Duan et al., 2022, Section 5.6). The four characters have appeared as CFT-type solutions of fourth order holomorphic MLDE in (Kaidi et al., 2021, Table 6), and their $q$-series expansions can be found in (Duan et al., 2022, Equation (5.34)). The $M_{2}$ can be realized by the coset construction $(D_{4})_{1}/(A_{2})_{1}$ (Duan et al., 2022, Section 5.6). We notice that it can also be realized as a non-diagonal modular invariant of $(U_{1})_{2}\times(U_{1})_{6}$, which is consistent with magic triangle element $U_{1}^{2}$.
@@
  • $\mathcal{T}^{(3/2,1)}=D_{2\rm A}$ is a unitary RCFT associated to the 2A conjugacy class of the Monster group. It has $c=6/5$ and conformal weights $h_{i}=0,(\frac{1}{10})_{3},(\frac{1}{2})_{3},\frac{3}{5}$. The corresponding VOA can be constructed by using a code over $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and as an extension of the product of Ising model $\mathrm{Vir}_{4,3}$ and tri-critical Ising model $\mathrm{Vir}_{5,4}$ Lam and Yamada (2000). For 2d RCFTs associated to various conjugacy classes and sporadic groups, see a good review in Bae et al. (2021b).
@@
  • $\mathcal{T}^{(3,2/3)}=\mathrm{Vir}_{6,5}^{\rm e}$ is a non-diagonal modular invariant of minimal model $\mathrm{Vir}_{6,5}$ which describes the critical 3-state Potts model. The corresponding VOA can be constructed by using a code over $\mathbb{Z}_{3}$ Miyamoto (2001).
@@
  • $\mathcal{T}^{(3/2,3/2)}=(A_{1})_{8}^{\rm e}$ is the D-type modular invariant of $(A_{1})_{8}$. The associated VOA can be constructed by simple current extension.
@@
All RCFTs associated with the extended magic triangle are presented in Table 2. The data of the extended magic triangle and the corresponding RCFTs are collected in the Tables 5, 6, 7 in the Appendix A.
@@
We now discuss the symmetry of the magic triangle at level one. It is straightforward to verify that under the exchange of $\mu$ and $\nu$, the central charge $c$, the conformal weight $h_{1}$ and the fourth order MLDE are invariant, while $h_{2},h_{3}$ are exchanged. This suggests that the $\mathcal{T}^{(\mu,\nu)}$ and $\mathcal{T}^{(\nu,\mu)}$ theories are isomorphic to each other, differing only by a relabeling of $\chi_{2}$ and $\chi_{3}$. Consequently, it suffices to present only half of the triangle, as shown in the Tables 5, 6, 7. In total, there are 30 non-equivalent RCFTs in the extended magic triangle at level one. The sixteen of these theories arise from ordinary WZW model of semisimple Lie groups, eight arise from intermediate VOAs, three arise from Virasoro minimal models, two arise from $U_{1}$ theories (that is, compact bosons on a circle), and the final theory $D_{2\rm A}$ arises from the product of Ising model and tri-critical Ising model and is associated with Monster group. For all 30 theories on the magic triangle at level one, the Fourier coefficients of the characters remain non-negative integers.
Several remarks are in order: